OpenStudy (anonymous):

4(x − 1) = 2(x + 2) + 6

OpenStudy (anonymous):

Distributive Property. a(b + c) = ab + ac So, distribute first... 4(x - 1) = 2(x + 2) + 6 4x - 4 = 2x + 4 + 6 Now, Combine like terms ~ 4x - 4 = 2x + 4 + 6 4x - 4 = 2x + 10 Add 4 to both sides. 4x = 2x + 14 Subtract 2x from both sides. 2x = 14 Divide by 2. x = 14 /2 x = 7 Hope this helps!

OpenStudy (anonymous):

@broweyebrit